OpenStudy (tyler444):

At the school store, colored pencils are reduced from $2.40 a box to $2.00 a box. What is the percent of decrease?

OpenStudy (tyler444):

please help will fan and medal

jhonyy9 (jhonyy9):

use the rule of 3 simple the difference between $240 and $2,00 is $0,40 - yes ? $2,40 a box ------------- 100% $0,40 decrease mean --- x% ------------------------------- x = 0,40*100/2,40 = 4/2,40 = ?

jhonyy9 (jhonyy9):

sorry there from $2,40 i missed the decimal point

jhonyy9 (jhonyy9):

do you understand the way ?

OpenStudy (tyler444):

no honestly

jhonyy9 (jhonyy9):

ok. so from 2,40 was decreased to 2,00 - yes ? so by how many dollars was decreased the price ?

OpenStudy (tyler444):

0.4

jhonyy9 (jhonyy9):

by the difference between 2,40 and 2,00 - yes ? yes you are right with 0,40

jhonyy9 (jhonyy9):

ok. so than you know that the initial prize was 2,40 dollars so this mean that in percent will mean 100% ok. ?

OpenStudy (tyler444):

yes

jhonyy9 (jhonyy9):

ok so now you need calculi how many percent mean this difference of 0,40 dollars - ok. ?

OpenStudy (tyler444):

like 20% ??

jhonyy9 (jhonyy9):

idk . bc. this you need calculi it using this rule of 3 simple what i ve wrote it above

jhonyy9 (jhonyy9):

2,40 mean 100% 0,40 mean x% ---------------- x = 0,40*100/2,40 = 40/2,40 = ?%

jhonyy9 (jhonyy9):

how many is 40 divide 2,40 ?

jhonyy9 (jhonyy9):

use a calculator

OpenStudy (tyler444):

16.6666667

OpenStudy (tyler444):

so 16 and 2/3%

jhonyy9 (jhonyy9):

yes sure right 16,6 % mean this $0,40 so this is the right correct answer sure hope you understand it now yes ?

OpenStudy (tyler444):

kinda

jhonyy9 (jhonyy9):

sorry what mean this kinda please ?

jhonyy9 (jhonyy9):

hope yes good luck